# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

### 19 Jan 2024 v2

@@ Abstract
Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes. Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high performance and rapid inference speeds in the domain of STR. Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder. This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR. Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition. The code is publicly available at https://github.com/cxfyxl/VIPTR.
@@ 1 Introduction
Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al. (2016); Guo et al. (2022); Yang et al. (2020); Guo et al. (2023); Zhou et al. (2023); Li et al. (2023a); Wu et al. (2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
@@
Scene text recognition initially relied on deep learning models like CRNNs Shi et al. (2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise. Recent researchers 2021 (2020); Zhang et al. (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss. From the perspective of simplifying the model architecture, a number of models used only CNN Borisyuk et al. (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Inspired by the pyramid structure of Swin Transformer Liu et al. (2021), a SVTR model Du et al. (2022) with single visual modal input only is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information. SVTR has four structural variants with different capacities.
@@
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for single visual modality. On one hand, it practices the effectiveness of replacing convolution with self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveal the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
@@
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al. (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic relationships between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al. (2021) combined with manhattan self-attention (MaSA) Fan et al. (2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al. (2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
@@ 2 Related Work
Figure 2: Overall architecture of the proposed VIPTR. It is a four-stage network with three-stage height progressively decreased. In each stage, a series of attention mixing blocks are carried out and followed by a subsampling or combining operation. In particular, VIPTR designs two visual feature fusion modes in vision permulation blocks: series and parallel. At last, the recognition is conducted by CTC decoder.
@@
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet He et al. (2016) introduced residual connections along with the batch normalization (BN) layer, which enabled training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model undergoes a transformation into a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
@@
ViT Dosovitskiy et al. (2020) is the first work to introduce Transformer Vaswani et al. (2017); Yang et al. (2022); Touvron et al. (2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT Touvron et al. (2021) introduces a new training strategy, enabling ViT to achieve superior performance compared to CNNs on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Currently, vision Transformers are integrated into STR task as visual feature encoders, leading to a significant improvement in accuracy. ViTSTR Atienza (2021), visionLAN Wang et al. (2021b), and MaskOCR Lyu et al. (2022) have all come close to the SOTA effect. Among them, SVTR-T is in the context of inputting constant-width text images, it both ensures high performance and creates the fastest recognition efficiency. Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens. As the resolution of the input image increases, the inference time consumption also increases quadratically. square growth.
@@
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al. (2017); Yang et al. (2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) Wang et al. (2021a) effectively extracts global information by utilizing sparse token-region relationships, and OSRA introduces overlapping spatial reduction (OSR) Lou et al. (2023) into SRA to achieve better performance by using larger and overlapping patches.
@@ 3.1 Design Guidelines
In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high performance networks for STR. Here are the design guidelines we followed:
@@ 3.2 Overall Architecture
The overview of the proposed VIPTR is illustrated in Figure 2. It is a four-stage network with three height progressively decreased layers dedicated to text recognition. For an image text of size $H\times W\times 3$, it is first transformed to $\frac{H}{4}\times\frac{W}{4}$ patches of channel dimension $C_{0}$ via a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Then, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension. Local and Global Attention blocks are used to focus on character feature extraction and dependency capture between character components. The backbone is used to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which perceives multi-granularity character features. Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
@@ 3.3 Patch Embedding
The first step of vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This schemeDu et al. (2022) gradually increases feature dimensions, which is beneficial to feature fusion.
@@ 3.4 Character and String Modeling
End-to-end text recognizers rely heavily on the mining of character features and string context. However, most of the existing studies use feature sequences with height $1$ to represent image text. Each feature corresponds to a thin image region, which is usually noisy, especially for irregular text. This is not the best way to describe the characters of an image. The vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds. This shows that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual relationships of strings.
More specifically, image text requires features to be extracted from two perspectives. The first is local character features, such as strokes and fonts. It encodes morphological features and correlations between different parts of a character. The second is inter-character context relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of mixing blocks to sense relevance by using self-attention with different receptive fields. Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features. When attention mixing block is input, the Layer Normalization (LN), Position Embedding (PE) are applied and then the self-attention mechanism is performed for dependency modeling. Next, LN and Feed Forward Network (FFN) are successively applied for feature fusion. Finally, the residual connection is used.
@@
The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3LABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
@@
As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual relationships between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, in order to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
Positional Embedding.
Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE)Vaswani et al. (2017), relative positional encoding (RPE) Liu et al. (2021) and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added onto the input feature. Locally-enhanced Positional Encoding (LePE)Dong et al. (2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
As mentioned above, the two types of hybrid blocks aim to extract complementary different features. In VIPTR, blocks are processed locally first and then globally. At each stage, Local and Global Attentions are recurrently applied multiple times for comprehensive feature extraction. Effect of the two kinds of blocks arrangements will be shown in next ablation experiments.

### 24 Jan 2024 v3

@@ Abstract
Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes. Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high-performance and rapid inference speeds in the domain of STR. Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder. This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR. Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition. The code is publicly available at https://github.com/cxfyxl/VIPTR.
@@ 1 Introduction
Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al. (2016); Guo et al. (2022); Yang et al. (2020); Guo et al. (2023); Li et al. (2023a); Wu et al. (2023); Zhou et al. (2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
@@
Scene text recognition initially relied on deep learning models like CRNNs Shi et al. (2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise. Recent researchers 2021 (2020); Zhang et al. (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss. From the perspective of simplifying the model architecture, some models used only CNN Borisyuk et al. (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Inspired by the pyramid structure of Swin Transformer Liu et al. (2021), the SVTR model Du et al. (2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information. SVTR has four structural variants with different capacities.
@@
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality. On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
@@
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al. (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al. (2021) combined with Manhattan self-attention (MaSA) Fan et al. (2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al. (2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
@@ 2 Related Work
Figure 2: Overall architecture of the proposed VIPTR. It is a four-stage network with three-stage height progressively decreasing. In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation. In particular, VIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel. At last, the recognition is conducted by the CTC decoder.
@@
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet He et al. (2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
@@
ViT Dosovitskiy et al. (2020) is the first work to introduce Transformer Vaswani et al. (2017); Yang et al. (2022); Touvron et al. (2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT Touvron et al. (2021) introduces a new training strategy, enabling ViT to achieve superior performance on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy. ViTSTR Atienza (2021), visionLAN Wang et al. (2021b), and MaskOCR Lyu et al. (2022) have all come close to the SOTA effect. SVTR-T both ensures high performance and creates the fastest recognition efficiency. Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens. As the resolution of the input image increases, the inference time consumption also increases quadratically.
@@
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al. (2017); Yang et al. (2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) Wang et al. (2021a) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) Lou et al. (2023) into SRA to achieve better performance by using larger and overlapping patches.
@@ 3.1 Design Guidelines
In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high-performance networks for STR. Here are the design guidelines we followed:
@@ 3.2 Overall Architecture
The schematic representation of the proposed VIPTR is depicted in Figure 2. It is a four-stage network with three height progressively decreased layers. Given an image text of size $H\times W\times 3$, it undergoes an initial transformation into $\frac{H}{4}\times\frac{W}{4}\times C_{0}$ patches through a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension. Both Local and Global Attention blocks are employed to concentrate on character feature extraction and capture dependencies between character components. The extractor serves to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which captures multi-granularity character features. Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
@@ 3.3 Patch Embedding
The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This scheme Du et al. (2022) gradually increases feature dimensions, which is beneficial to feature fusion.
@@ 3.4 Character and String Modeling
Existing studies mainly use feature sequences with height $1$ to represent image text. Each feature corresponds to a small local region. However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well. In contrast, the Vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds. This suggests that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual dependencies of strings.
More specifically, image text requires features to be extracted from two perspectives. The first involves local character features, including elements like strokes and fonts, representing internal associations of a character and morphological features. The second is context relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of attention-mixing blocks with different receptive fields to sense relevance. Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features. Upon the input of the attention mixing block, Layer Normalization (LN) and Position Embedding (PE) are first applied. Subsequently, the self-attention mechanism is employed for dependency modeling. Finally, LN and Feed Forward Network (FFN) are successively applied for feature fusion, with the residual connection applied.
Positional Embedding.
Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE) Vaswani et al. (2017), relative positional encoding (RPE) Liu et al. (2021), and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added to the input feature. Locally-enhanced Positional Encoding (LePE) Dong et al. (2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
@@
The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3LABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than the Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing, and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
@@
As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
As mentioned above, the hybrid blocks in VIPTR are designed to extract distinct yet complementary features. In this way, blocks are initially processed locally before being addressed globally. At each stage, Local and Global Attention are recurrently applied multiple times for comprehensive feature extraction. Effect of the two kinds of blocks Permutations can be seen in ablation experiments.